OpenStudy (anonymous):

Object A travels in the +x-direction before hitting a stationary object B. Afterwards, object A’s x-momentum is 5.7 × 10^4 kilogram meters/second and its y-momentum is 6.2 × 10^4 kilogram meters/second. What is the object’s resultant angle of motion with the +x-axis after the collision?

OpenStudy (anonymous):

Any ideas?

OpenStudy (anonymous):

\[\[\tan \theta =\frac{ 6.2 \times 10^{4} }{ 5.7 \times 10 ^{4} }\]\]